OpenStudy (anonymous):

If f(x) = (5x + 3)/(4x-5), x is not equal to 5/4, show that fof is an identity function.

OpenStudy (hba):

WELCOME TO OPEN STUDY

OpenStudy (anonymous):

thank yu

OpenStudy (vishweshshrimali5):

:) Welcome @ashna

OpenStudy (vishweshshrimali5):

Do you know what is the meaning of an identity function and fof.......

OpenStudy (anonymous):

\[ f(f(x))=\frac{\frac{5 (5 x+3)}{4 x-5}+3}{\frac{4 (5 x+3)}{4 x-5}-5} \]

OpenStudy (vishweshshrimali5):

Now solve this................

OpenStudy (anonymous):

no i dont know .. @eliassaab should i solve that ? .. is the answer x ?

OpenStudy (anonymous):

\[ f(f(x))=\frac{\frac{5 (5 x+3)}{4 x-5}+3}{\frac{4 (5 x+3)}{4 x-5}-5} =\frac{37 x}{\frac{37 (4 x-5)}{4 x-5}}=x \]

OpenStudy (vishweshshrimali5):

Identity function : a function which is true for all values of the variable. and fof means function f of f(x)

OpenStudy (anonymous):

oh .. thanks :)

OpenStudy (anonymous):

Did you understand it?

OpenStudy (vishweshshrimali5):

Now as u get the value x it means that the value of fof is a real value for any value of x except 5/4............ so fof becomes an identity function

OpenStudy (anonymous):

yes .. thank you .. can i ask another question ?

OpenStudy (vishweshshrimali5):

@ashna Ask it as a seperate question

OpenStudy (anonymous):

okay !

OpenStudy (hba):

u would have to close this question to post another @ashna